A metal PQ of mass 10g lies at rest on two horizontal metal rails seperated by 4.90cm.A vertically downward magnetic feild of 0.800T exists in the space.the resistance of the circuit is slowly decreased and it is found that when the resistance goes below 20 omega,the wire PQ starts sliding on the rails.Find the coefficient of friction.

  1. A metal PQ of mass 10g lies at rest on two horizontal metal rails seperated by 4.90cm.A vertically downward magnetic feild of 0.800T exists in the space.the resistance of the circuit is slowly decreased and it is found that when the resistance goes below 20 omega,the wire PQ starts sliding on the rails.Find the coefficient of friction.

Grade:12

1 Answers

ROSHAN MUJEEB
3 years ago
Solution :
μR=F
rArrμ×m×g=iLB
μ×0.10×10−3×9.8
=620×4.9×10−2×0.8
μ=6×4.9×10−2×0.820×0.10×10−2×9.8
μ=0.12 .